# Kineo: Calibration-Free Metric Motion Capture From Sparse RGB Cameras

### v1

@@ 5.5 Optimal Absolute Extrinsics
$R\mid\mathbf{t}]_{k}^{*}=\penalty 10000\ \prod_{(i,j)\penalty 10000\ \in P^{*}_{k}}^{\leftarrow}[R\mid\mathbf{t}]_{ij}=\prod_{(i,j)\penalty 10000\ \in P^{*}_{k}}^{\leftarrow}\penalty 10000\ \begin{bmatrix}R_{ij}&\penalty 10000\ \lambda_{ij}\mathbf{\hat{t_{ij}}}\penalty 10000\ \\ \penalty 10000\ \mathbf{0}^{T}&1\end{bmatrix}$, (7)
@@ 5.6 Bundle Adjustment
$\sum_{i=1}^{N}\mathbf{1}\!\left[w_{i,k}>\kappa\right]\;\geq\;2,$ (8)
@@
$\epsilon_{i,m}=\|\pi(K_{i},\mathbf{d}_{i},[R|\mathbf{t}]_{i},\mathbf{X}_{m})-\mathbf{x}_{m}\|_{2}$ (9)
@@
$\mathcal{L}_{\text{reproj}}=\frac{\sum_{i=1}^{N}\sum_{m=1}^{M}w_{i,m}\mathcal{H}_{\delta}\left(\epsilon_{i,m}\right)}{\sum_{i=1}^{N}\sum_{m=1}^{M}w_{i,m}}$ (10)
@@
6 3D Output And Confidence Score
@@
6.1 Scene Pointmaps
We evaluated two models to reconstruct the scene pointmaps: MoGe [54] and VGGT [53]. For both methods, each input image is first undistorted using the predicted distortion coefficients.
@@ 6.1 Scene Pointmaps
Figure 4: Comparison of scene pointmaps produced by two geometry estimation models. (a) shows the pointmap generated with MoGe, which uses single-view images and camera intrinsics to produce metric pointmaps. (b) shows the pointmap generated with VGGT, which leverages multiple views to predict a globally consistent pointmap and camera poses, aligned to our camera system. VGGT provides more spatially consistent pointmaps across views, while MoGe relies on single-view depth estimation.
6.2 Triangulation and Confidence Score
The 2D keypoints are lifted to 3D space through a triangulation process using a weighted Direct Linear Transform (DLT) method. This process begins by first undistorting the 2D keypoints using the estimated distortion parameters. Subsequently, the undistorted keypoints are triangulated to determine their 3D positions, leveraging information from all N views. Crucially, the triangulation is weighted by the 2D confidence score of each individual keypoint, $w_{i,m}$. This weighting allows the reconstruction to account for varying levels of uncertainty across different views, thereby improving the accuracy and robustness of the final 3D keypoint coordinates. As the pipeline takes any 2D keypoints as input, the system can generate multiple output formats to suit diverse downstream applications. Figure 5 illustrates this flexibility, showing examples of 3D reconstructions in COCO-19, COCO-133, Human3.6M keypoint formats, as well as a full parametric SMPL body model. The SMPL reconstruction is obtained by first detecting NLF [39] surface keypoints and subsequently fitting them using the SMPLFitter module from NLF.
@@ 6.2 Triangulation and Confidence Score
Figure 5: Examples of different output formats with different 2D keypoints detectors used by Kineo. (a–c) show keypoint-based representations in various formats, while (d) illustrates a full parametric SMPL body reconstruction. For the SMPL model, NLF surface keypoints are first detected and subsequently fitted using the SMPLFitter module from NLF.
6.3 Pairwise Reprojection Confidence Score
@@
$\tilde{\epsilon}_{i,m}=\frac{\|\pi(K_{i},\mathbf{d}_{i},[R|\mathbf{t}]_{i},\mathbf{X}_{m})-\mathbf{x}_{m}\|_{2}}{f_{i}}$ (11)
@@
$s_{i,m}=\exp(-\lambda\tilde{\epsilon}_{i,m})$ (12)
@@
$s_{ij,m}=\sqrt{w_{i}\cdot w_{j}}\sqrt{s_{i,m}\cdot s_{j,m}}$ (13)
@@
Figure 6: Visualization of the proposed pairwise reprojection concensus score on 3D human skeleton reconstructions from the EgoHumans dataset, obtained using Kineo. The low score in poorly reconstructed hands demonstrates how the proposed confidence correlates with reconstruction quality and could be leveraged for downstream filtering; note that we did not implement this filtering in our method and leave it for future work.
7 Metric Scale Recovery
@@
7.1 SMPL-based Scale Estimation
Figure 7: Overview of SMPL-based global scaling. The SMPL body model serves as a prior for metric scale, providing plausible human body proportions that guide the estimation of the global scale factor $\alpha$, which converts the scene from arbitrary units to metric units. Given an estimated skeleton in arbitrary units, we optimize $\alpha$ while updating the SMPL shape parameters $\beta$ so that the SMPL skeleton closely matches the proportions of the input skeleton, enforced by $\mathcal{L}_{\text{bones}}$ while remaining in the space of plausible, metric-consistent shapes, enforced by $\mathcal{L}_{\beta}$. The resulting optimized scale $\alpha^{*}$ is then applied to rescale the entire scene, including the skeletons, cameras, and scene pointmaps, producing a metric-scale representation.
The first approach estimates the global scale factor by leveraging the SMPL parametric body model. Conceptually, this approach is related to the method of Müller et al. [29], which incorporates the SMPL model within a global optimization loop to constrain the solution space to metric-consistent configurations. However, our method differs in how the SMPL body prior is used to extract the global scale. Specifically, we treat scale estimation as a downstream task rather than integrating it into a global optimization. This avoids potential conflicts between camera and body shape objectives, reducing the risk of being stucked in local minima. Given a skeleton in arbitrary units, we formulate an optimization problem to recover the correct global scale. The goal is to determine the scale factor $\alpha$ that maps the arbitrary-unit skeleton to metric space while enforcing the same body proportions between the two skeletons by updating the SMPL shape parameters $\beta$. This optimization is computationally efficient, as it only requires estimating body shape, ignoring pose, and thus eliminating the need for linear blend skinning. This approach is illustrated in Fig. 7.
@@ 7.1 SMPL-based Scale Estimation
$\hat{\mathbf{j}}(\beta)=M(\beta)\cdot J$ (15)
@@
$\ell_{kl}=\alpha\|\mathbf{j}_{k}-\mathbf{j}_{l}\|_{2}\quad\text{and}\quad\hat{\ell}_{kl}=\|\hat{\mathbf{j}}_{k}-\hat{\mathbf{j}}_{l}\|_{2}\quad(k,l)\in S$ (16)
@@
$\beta_{0}=\mathbf{0}\quad\text{and}\quad\alpha_{0}=\frac{\sum_{(k,l)\in S}s_{kl}\,\frac{\hat{\ell}_{kl}}{\ell_{kl}}}{\sum_{(k,l)\in S}s_{kl}}$ (18)
@@
$\mathcal{L}_{\text{bones}}=\frac{\sum_{(k,l)\in S}s_{kl}\,\mathcal{H}_{\delta}(\ell-\hat{\ell})}{\sum_{(k,l)\in S}s_{kl}}\quad\text{and}\quad\mathcal{L}_{\beta}=\|\beta\|_{2}^{2}$ (19)
@@
$\mathcal{L}_{\text{tot}}=\lambda_{\text{bones}}\mathcal{L}_{\text{bones}}+\lambda_{\beta}\mathcal{L}_{\beta}$ (20)
@@
7.2 Metric Depth-based Scale Estimation
Figure 8: Overview of metric depth-based global scaling. Unlike SMPL-based scaling, which relies on a parametric human body model, metric depth-based scaling is subject-agnostic, as the metric depth map predicted directly from a the RGB image by MoGe [54] is independent of the subjects present in the scene. The global scale factor $\alpha$ is computed as the ratio between the joint depths in metric units obtained from the depth map and the corresponding joint depths in arbitrary units. The resulting scale $\alpha$ is then applied to rescale the entire scene, including the skeletons, cameras, and scene pointmaps, producing a metric-scale representation.
While the previous SMPL-based approach is effective at recovering global scale, it is limited to footage including humans due to its reliance on a parametric body model. We overcome this restriction by using a monocular metric depth estimator, MoGe [54], which can estimate a metric depth map from a monocular RGB image. Since MoGe can take the focal length as input to improve output quality, we feed it the intrinsics estimated in previous stages (see Sec. 5.1 and Sec. 5.6). The global scale $\alpha$ is computed as the ratio between the joint depths in metric units obtained from the depth map and the corresponding joint depths in arbitrary units. This approach is illustrated in Fig. 8.
@@ 7.2 Metric Depth-based Scale Estimation
$\mathbf{j}_{\text{cam}}=\begin{bmatrix}x_{c}\\ y_{c}\\ z_{c}\end{bmatrix}=[R\mid\mathbf{t}]\cdot\mathbf{j},\quad d_{a}=z_{c}$ (21)
@@
$\mathbf{j}_{\text{img}}=\begin{bmatrix}u\\ v\end{bmatrix}=\pi(K,\mathbf{d},[R|\mathbf{t}],\mathbf{j}),\quad d_{m}=D_{v,u}$ (22)
@@
$h_{i}=\frac{1}{K}\sum^{K}_{k=1}\exp(-\lambda\epsilon_{i,k})$ (23)
@@
$\alpha=\frac{\sum_{i=1}^{N}\sum_{k=1}^{K}w_{i,k},\frac{d_{m,k}}{d_{a,k}}}{\sum_{i=1}^{N}\sum_{k=1}^{K}w_{i,k}}.$ (24)
@@ 8.1 Experimental Setup
We evaluate our method on two complementary datasets that together provide comprehensive coverage of both controlled and in-the-wild scenarios. Human3.6M [16] provides highly accurate ground-truth annotations from a Vicon motion capture system, making it well-suited for evaluating joint localization accuracy and inference speed in controlled, single-person scenarios. EgoHumans [19] presents more complex scenarios with both indoor and outdoor settings, multiple interacting subjects, frequent occlusions, and dynamic motions, enabling a thorough assessment of our method’s robustness under realistic, unconstrained conditions. Notably, EgoHumans recordings are captured with GoPro cameras that exhibit significant lens distortion, making this dataset particularly valuable for evaluating the quality of our distortion coefficient estimation, a capability that distinguishes our method from prior approaches that typically ignore distortions or undistort the images with the ground-truth distortions coefficients.
@@
We assess both joint position accuracy and camera parameters quality using a comprehensive set of standard metrics. Following Müller et al. [29], the predicted scene is first aligned with the ground truth based on camera centers prior to evaluation. For 3D joints reconstruction, we report the World Mean Per Joint Position Error (W-MPJPE), which quantifies the average Euclidean distance between predicted and ground-truth joint positions in the metric world coordinate system. To isolate local pose errors from global misalignment and scale variations, we also report the Procrustes-Aligned MPJPE (PA-MPJPE), computed per subject and per frame after applying a similarity transform from the predicted skeleton to the ground truth. For camera pose, we report the Translation Error (TE), in meters, and Angular Error (AE), in degrees, to quantify deviations in position and orientation from the ground truth. Complementing these error metrics, we report the Camera Center Accuracy (CCA) [26], which measures the proportion of estimated camera centers within predefined distance thresholds relative to the scene scale, and the Relative Rotation Accuracy (RRA), defined as the percentage of cameras whose angular error falls below specified thresholds. To decouple positional errors from global scale, we additionally compute scale-aligned variants by applying a similarity transformation that aligns and uniformly scales the estimated scene to the ground truth, resulting in the scale-aligned Translation Error (s-TE) and scale-aligned Camera Center Accuracy (s-CCA). Finally, we report the Field-of-View error (FoV), in degrees, to quantify deviations between the estimated and ground-truth camera field of view.
@@
8.2 Quantitative Evaluation
We benchmark our method against state-of-the-art multi-view reconstruction approaches that utilize different sets of ground truth camera parameters. UnCaliPose [56] uses known camera intrinsics while estimating extrinsics, whereas HAMSt3R [38] and HSfM [29] jointly estimate both intrinsics and extrinsics, with HSfM representing the current state-of-the-art calibration-free method on EgoHumans. We also compare to methods evaluated with both known intrinsics and extrinsics on Human3.6M, including AdaFuse [62] and Iskakov et al. [17], as well as Hewitt et al. [14], which, although presented as calibration-free, relies on both parameters in this benchmark. Our evaluation considers multiple configurations in which camera parameters are either estimated, provided as ground truth. We also assess performance both with and without modeling lens distortions to quantify their impact on reconstruction accuracy. Tables 1 and 2 report the human reconstruction and camera metrics, respectively. We evaluate HSfM accuracy on Human3.6M using the official code released on the authors’ repository, ensuring evaluation conditions consistent with those used for EgoHumans and EgoExo4D. For full transparency, inference scripts are available at https://github.com/cjaverliat/hsfm, and the evaluation code is included in Kineo’s repository.

### v2

@@ 5.5. Optimal Absolute Extrinsics
(7) $R\mid\mathbf{t}]_{k}^{*}=~\prod_{(i,j)~\in P^{*}_{k}}^{\leftarrow}[R\mid\mathbf{t}]_{ij}=\prod_{(i,j)~\in P^{*}_{k}}^{\leftarrow}~\begin{bmatrix}R_{ij}&~\lambda_{ij}\mathbf{\hat{t_{ij}}}~\\ ~\mathbf{0}^{T}&1\end{bmatrix}$,
@@ 5.6. Bundle Adjustment
(8) $\sum_{i=1}^{N}\mathbf{1}\!\left[w_{i,k}>\kappa\right]\;\geq\;2,$
@@
(9) $\epsilon_{i,m}=\|\pi(K_{i},\mathbf{d}_{i},[R|\mathbf{t}]_{i},\mathbf{X}_{m})-\mathbf{x}_{m}\|_{2}$
@@
(10) $\mathcal{L}_{\text{reproj}}=\frac{\sum_{i=1}^{N}\sum_{m=1}^{M}w_{i,m}\mathcal{H}_{\delta}\left(\epsilon_{i,m}\right)}{\sum_{i=1}^{N}\sum_{m=1}^{M}w_{i,m}}$
@@
6. 3D Output And Confidence Score
@@
6.1. Scene Pointmaps
We evaluated two models to reconstruct the scene pointmaps: MoGe (Wang et al., 2025b) and VGGT (Wang et al., 2025a). For both methods, each input image is first undistorted using the predicted distortion coefficients.
@@ 6.1. Scene Pointmaps
Figure 4. Comparison of scene pointmaps produced by two geometry estimation models. (a) shows the pointmap generated with MoGe, which uses single-view images and camera intrinsics to produce metric pointmaps. (b) shows the pointmap generated with VGGT, which leverages multiple views to predict a globally consistent pointmap and camera poses, aligned to our camera system. VGGT provides more spatially consistent pointmaps across views, while MoGe relies on single-view depth estimation.
6.2. Triangulation and Confidence Score
The 2D keypoints are lifted to 3D space through a triangulation process using a weighted Direct Linear Transform (DLT) method. This process begins by first undistorting the 2D keypoints using the estimated distortion parameters. Subsequently, the undistorted keypoints are triangulated to determine their 3D positions, leveraging information from all N views. Crucially, the triangulation is weighted by the 2D confidence score of each individual keypoint, $w_{i,m}$. This weighting allows the reconstruction to account for varying levels of uncertainty across different views, thereby improving the accuracy and robustness of the final 3D keypoint coordinates. As the pipeline takes any 2D keypoints as input, the system can generate multiple output formats to suit diverse downstream applications. Figure 5 illustrates this flexibility, showing examples of 3D reconstructions in COCO-19, COCO-133, Human3.6M keypoint formats, as well as a full parametric SMPL body model. The SMPL reconstruction is obtained by first detecting NLF (Sárándi and Pons-Moll, 2025) surface keypoints and subsequently fitting them using the SMPLFitter module from NLF.
@@ 6.2. Triangulation and Confidence Score
Figure 5. Examples of different output formats with different 2D keypoints detectors used by Kineo. (a–c) show keypoint-based representations in various formats, while (d) illustrates a full parametric SMPL body reconstruction. For the SMPL model, NLF surface keypoints are first detected and subsequently fitted using the SMPLFitter module from NLF.
6.3. Pairwise Reprojection Confidence Score
@@
(11) $\tilde{\epsilon}_{i,m}=\frac{\|\pi(K_{i},\mathbf{d}_{i},[R|\mathbf{t}]_{i},\mathbf{X}_{m})-\mathbf{x}_{m}\|_{2}}{f_{i}}$
@@
(12) $s_{i,m}=\exp(-\lambda\tilde{\epsilon}_{i,m})$
@@
(13) $s_{ij,m}=\sqrt{w_{i}\cdot w_{j}}\sqrt{s_{i,m}\cdot s_{j,m}}$
@@
Figure 6. Visualization of the proposed pairwise reprojection concensus score on 3D human skeleton reconstructions from the EgoHumans dataset, obtained using Kineo. The low score in poorly reconstructed hands demonstrates how the proposed confidence correlates with reconstruction quality and could be leveraged for downstream filtering; note that we did not implement this filtering in our method and leave it for future work.
7. Metric Scale Recovery
@@
7.1. SMPL-based Scale Estimation
Figure 7. Overview of SMPL-based global scaling. The SMPL body model serves as a prior for metric scale, providing plausible human body proportions that guide the estimation of the global scale factor $\alpha$, which converts the scene from arbitrary units to metric units. Given an estimated skeleton in arbitrary units, we optimize $\alpha$ while updating the SMPL shape parameters $\beta$ so that the SMPL skeleton closely matches the proportions of the input skeleton, enforced by $\mathcal{L}_{\text{bones}}$ while remaining in the space of plausible, metric-consistent shapes, enforced by $\mathcal{L}_{\beta}$. The resulting optimized scale $\alpha^{*}$ is then applied to rescale the entire scene, including the skeletons, cameras, and scene pointmaps, producing a metric-scale representation.
The first approach estimates the global scale factor by leveraging the SMPL parametric body model. Conceptually, this approach is related to the method of Müller et al. (Müller et al., 2025), which incorporates the SMPL model within a global optimization loop to constrain the solution space to metric-consistent configurations. However, our method differs in how the SMPL body prior is used to extract the global scale. Specifically, we treat scale estimation as a downstream task rather than integrating it into a global optimization. This avoids potential conflicts between camera and body shape objectives, reducing the risk of being stucked in local minima. Given a skeleton in arbitrary units, we formulate an optimization problem to recover the correct global scale. The goal is to determine the scale factor $\alpha$ that maps the arbitrary-unit skeleton to metric space while enforcing the same body proportions between the two skeletons by updating the SMPL shape parameters $\beta$. This optimization is computationally efficient, as it only requires estimating body shape, ignoring pose, and thus eliminating the need for linear blend skinning. This approach is illustrated in Fig. 7.
@@ 7.1. SMPL-based Scale Estimation
(15) $\hat{\mathbf{j}}(\beta)=M(\beta)\cdot J$
@@
(16) $\ell_{kl}=\alpha\|\mathbf{j}_{k}-\mathbf{j}_{l}\|_{2}\quad\text{and}\quad\hat{\ell}_{kl}=\|\hat{\mathbf{j}}_{k}-\hat{\mathbf{j}}_{l}\|_{2}\quad(k,l)\in S$
@@
(18) $\beta_{0}=\mathbf{0}\quad\text{and}\quad\alpha_{0}=\frac{\sum_{(k,l)\in S}s_{kl}\,\frac{\hat{\ell}_{kl}}{\ell_{kl}}}{\sum_{(k,l)\in S}s_{kl}}$
@@
(19) $\mathcal{L}_{\text{bones}}=\frac{\sum_{(k,l)\in S}s_{kl}\,\mathcal{H}_{\delta}(\ell-\hat{\ell})}{\sum_{(k,l)\in S}s_{kl}}\quad\text{and}\quad\mathcal{L}_{\beta}=\|\beta\|_{2}^{2}$
@@
(20) $\mathcal{L}_{\text{tot}}=\lambda_{\text{bones}}\mathcal{L}_{\text{bones}}+\lambda_{\beta}\mathcal{L}_{\beta}$
@@
7.2. Metric Depth-based Scale Estimation
Figure 8. Overview of metric depth-based global scaling. Unlike SMPL-based scaling, which relies on a parametric human body model, metric depth-based scaling is subject-agnostic, as the metric depth map predicted directly from a the RGB image by MoGe (Wang et al., 2025b) is independent of the subjects present in the scene. The global scale factor $\alpha$ is computed as the ratio between the joint depths in metric units obtained from the depth map and the corresponding joint depths in arbitrary units. The resulting scale $\alpha$ is then applied to rescale the entire scene, including the skeletons, cameras, and scene pointmaps, producing a metric-scale representation.
While the previous SMPL-based approach is effective at recovering global scale, it is limited to footage including humans due to its reliance on a parametric body model. We overcome this restriction by using a monocular metric depth estimator, MoGe (Wang et al., 2025b), which can estimate a metric depth map from a monocular RGB image. Since MoGe can take the focal length as input to improve output quality, we feed it the intrinsics estimated in previous stages (see Sec. 5.1 and Sec. 5.6). The global scale $\alpha$ is computed as the ratio between the joint depths in metric units obtained from the depth map and the corresponding joint depths in arbitrary units. This approach is illustrated in Fig. 8.
@@ 7.2. Metric Depth-based Scale Estimation
(21) $\mathbf{j}_{\text{cam}}=\begin{bmatrix}x_{c}\\ y_{c}\\ z_{c}\end{bmatrix}=[R\mid\mathbf{t}]\cdot\mathbf{j},\quad d_{a}=z_{c}$
@@
(22) $\mathbf{j}_{\text{img}}=\begin{bmatrix}u\\ v\end{bmatrix}=\pi(K,\mathbf{d},[R|\mathbf{t}],\mathbf{j}),\quad d_{m}=D_{v,u}$
@@
(23) $h_{i}=\frac{1}{K}\sum^{K}_{k=1}\exp(-\lambda\epsilon_{i,k})$
@@
(24) $\alpha=\frac{\sum_{i=1}^{N}\sum_{k=1}^{K}w_{i,k},\frac{d_{m,k}}{d_{a,k}}}{\sum_{i=1}^{N}\sum_{k=1}^{K}w_{i,k}}.$
@@ 8.1. Experimental Setup
We evaluate our method on two complementary datasets that together provide comprehensive coverage of both controlled and in-the-wild scenarios. Human3.6M (Ionescu et al., 2014) provides highly accurate ground-truth annotations from a Vicon motion capture system, making it well-suited for evaluating joint localization accuracy and inference speed in controlled, single-person scenarios. EgoHumans (Khirodkar et al., 2023) presents more complex scenarios with both indoor and outdoor settings, multiple interacting subjects, frequent occlusions, and dynamic motions, enabling a thorough assessment of our method’s robustness under realistic, unconstrained conditions. Notably, EgoHumans recordings are captured with GoPro cameras that exhibit significant lens distortion, making this dataset particularly valuable for evaluating the quality of our distortion coefficient estimation, a capability that distinguishes our method from prior approaches that typically ignore distortions or undistort the images with the ground-truth distortions coefficients.
@@
We assess both joint position accuracy and camera parameters quality using a comprehensive set of standard metrics. Following Müller et al. (Müller et al., 2025), the predicted scene is first aligned with the ground truth based on camera centers prior to evaluation. For 3D joints reconstruction, we report the World Mean Per Joint Position Error (W-MPJPE), which quantifies the average Euclidean distance between predicted and ground-truth joint positions in the metric world coordinate system. To isolate local pose errors from global misalignment and scale variations, we also report the Procrustes-Aligned MPJPE (PA-MPJPE), computed per subject and per frame after applying a similarity transform from the predicted skeleton to the ground truth. For camera pose, we report the Translation Error (TE), in meters, and Angular Error (AE), in degrees, to quantify deviations in position and orientation from the ground truth. Complementing these error metrics, we report the Camera Center Accuracy (CCA) (Lin et al., 2024), which measures the proportion of estimated camera centers within predefined distance thresholds relative to the scene scale, and the Relative Rotation Accuracy (RRA), defined as the percentage of cameras whose angular error falls below specified thresholds. To decouple positional errors from global scale, we additionally compute scale-aligned variants by applying a similarity transformation that aligns and uniformly scales the estimated scene to the ground truth, resulting in the scale-aligned Translation Error (s-TE) and scale-aligned Camera Center Accuracy (s-CCA). Finally, we report the Field-of-View error (FoV), in degrees, to quantify deviations between the estimated and ground-truth camera field of view.
@@
8.2. Quantitative Evaluation
We benchmark our method against state-of-the-art multi-view reconstruction approaches that utilize different sets of ground truth camera parameters. UnCaliPose (Xu and Kitani, 2022) uses known camera intrinsics while estimating extrinsics, whereas HAMSt3R (Rojas et al., 2025) and HSfM (Müller et al., 2025) jointly estimate both intrinsics and extrinsics, with HSfM representing the current state-of-the-art calibration-free method on EgoHumans. We also compare to methods evaluated with both known intrinsics and extrinsics on Human3.6M, including AdaFuse (Zhang et al., 2021) and Iskakov et al. (Iskakov et al., 2019), as well as Hewitt et al. (Hewitt et al., 2024b), which, although presented as calibration-free, relies on both parameters in this benchmark. Our evaluation considers multiple configurations in which camera parameters are either estimated, provided as ground truth. We also assess performance both with and without modeling lens distortions to quantify their impact on reconstruction accuracy. Tables 1 and 2 report the human reconstruction and camera metrics, respectively. We evaluate HSfM accuracy on Human3.6M using the official code released on the authors’ repository, ensuring evaluation conditions consistent with those used for EgoHumans and EgoExo4D. For full transparency, inference scripts are available at https://github.com/cjaverliat/hsfm, and the evaluation code is included in Kineo’s repository.